OpenStudy (anonymous):

Part A: Use the properties of exponents to explain why 64 raised to the power of 1 over 2 is called the square root of 64. Part B: The length of a rectangle is 7 units and its width is square root of 7 unit. Is the area of the rectangle rational or irrational?

OpenStudy (anonymous):

hint: \(x^a\cdot x^b = x^{a+b}\)

OpenStudy (anonymous):

for part A.

OpenStudy (anonymous):

okay so how would i put the answer

OpenStudy (anonymous):

use the property...

OpenStudy (anonymous):

if a = 1/2 and b = 1/2 then what's \(\huge{64^{\frac{1}{2}}\cdot64^{\frac{1}{2}}}\)

OpenStudy (anonymous):

is it becouse 8*8 = 64 and 64 1/2 is square root of 64

OpenStudy (anonymous):

\[64^{\frac{1}{2}}\cdot64^{\frac{1}{2}}=64^1\] so what times itself = 64? \[\sqrt{64}\cdot\sqrt{64} = 64\Rightarrow \sqrt{64}=64^{\frac{1}{2}}\]

OpenStudy (anonymous):

also\[\left( 64^{\frac{1}{2}} \right)^2 =64^{\frac{1}{2}\cdot2}=64^{1} \Rightarrow 64^{\frac{1}{2}}=\sqrt{64}\]

OpenStudy (anonymous):

8

OpenStudy (anonymous):

just using a different property of exponents

OpenStudy (anonymous):

it's not about \(64^{\frac{1}{2}}=8\) but \(x^{\frac{1}{2}}=\sqrt{x}\)

OpenStudy (anonymous):

oh okay i get it now!

OpenStudy (anonymous):

good!

OpenStudy (anonymous):

can u help me with part b

OpenStudy (anonymous):

what's the area?

OpenStudy (anonymous):

would it be 14

OpenStudy (anonymous):

length times width

OpenStudy (anonymous):

42

OpenStudy (anonymous):

what's the length?

OpenStudy (anonymous):

7

OpenStudy (anonymous):

and what's the width?

OpenStudy (anonymous):

7 squared

OpenStudy (anonymous):

square root of 7

OpenStudy (anonymous):

there you go. so multiply them \[7\cdot \sqrt{7}=?\]

OpenStudy (anonymous):

18.2?

OpenStudy (anonymous):

\[7\sqrt{7}\]

OpenStudy (anonymous):

is that rational or irrational?

OpenStudy (anonymous):

rational

OpenStudy (anonymous):

really? is \(\sqrt{7}\) rational?

OpenStudy (anonymous):

irrational? 0_o i suck at math.

OpenStudy (anonymous):

don't say that! you can say you're working to improve. keep it positive! rational simply means that a number can be expressed as a fraction. \(\sqrt{7}\) cannot be expressed as a fraction. it can be approximated with a fraction but there is no fraction involving only integers which equals \(\sqrt{7}\).

OpenStudy (anonymous):

fraction = ratio of integers

OpenStudy (anonymous):

or actually, rational number means "ratio of integers."

OpenStudy (anonymous):

:O thanks! u helped a lot!

OpenStudy (anonymous):

you're welcome!